# Incompressible Limit of Compressible Ideal MHD Flows inside a Perfectly Conducting Wall

### v4

@@ 3.2.2 Div-Curl analysis
The term $\|\varepsilon^{4+(k-1)_{+}}\partial_{t}^{k}\mathcal{T}^{2}(D_{t}p)\|_{2-k}^{2}$ will be further reduced to the tangenital estimates of $(u,B)$ by using the argument in Section 3.2.1. Finally, we need to control the tangential derivatives of $u$ and $B$ (including time derivatives) in $E_{5}(t)$ and the $L^{2}$ norm of $\partial_{t}^{5}p$ with suitable Mach number weights. Next, we analyze the vorticity term. The proof is parallel to Lemma 3.2.
@@
$\sum_{k=0}^{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}u\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}+1}B\times\left(\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B\right)\right\|_{2-k}^{2}\right)$ (3.29)
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho\left|\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}u\right|^{2}+\left|\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\right|^{2}+\left|\varepsilon^{3}B\times(\partial^{2}\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t),E_{5}(t),E_{6}(t)).$
@@
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}|\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B|^{2}\,\mathrm{d}x\underbrace{+\int_{\Omega}\left(\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\right)\cdot\left(\varepsilon^{4}\partial^{2}\nabla\times\mathcal{T}^{2}(BD_{t}p)\right)\,\mathrm{d}x}_{:=\mathcal{K}^{\prime}}$
$\underbrace{-\int_{\Omega}\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\cdot\varepsilon^{2}\left([\partial^{2}\nabla\times\mathcal{T}^{2},D_{t}]B+(u\cdot\nabla)(\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B)+[B\cdot\nabla,\partial^{2}\nabla\times\mathcal{T}^{2}]u\right)\,\mathrm{d}x}_{:=\mathcal{I}_{3}^{\prime}}$ (3.32)
@@
$\varepsilon^{2}\|\mathcal{T}(\rho u_{j})\partial^{3}\mathcal{T}\partial_{j}u_{i}\|_{0}\leq$ $~{}\|\mathcal{T}(\rho u)\|_{\infty}\|\varepsilon^{2}\mathcal{T}\partial_{j}u\|_{3}\leq P(E_{4}(t),E_{5}(t)),~{}j=1,2,$ (3.36)
@@
For $\mathcal{K}^{\prime}$, we consider the estimate of $\varepsilon^{4}\partial^{2}\nabla\times\mathcal{T}^{2}(BD_{t}p)$. Since,
$\varepsilon^{4}\partial^{2}\nabla\times\mathcal{T}^{2}(BD_{t}p)$
$=$ $~{}\varepsilon^{4}\partial^{2}\mathcal{T}^{2}(D_{t}\nabla p\times B)\underbrace{+\varepsilon^{4}\partial^{2}\mathcal{T}^{2}(\nabla u\cdot\nabla p\times B)+\varepsilon^{4}\partial^{2}\mathcal{T}^{2}(D_{t}p\nabla\times B)}_{:=\mathcal{R}_{1}^{\prime}}$
$=$ $-B\times\varepsilon^{4}\partial^{2}\mathcal{T}^{2}D_{t}\nabla p\underbrace{-\varepsilon^{4}[\partial^{2}\mathcal{T}^{2},B\times](D_{t}\nabla p)}_{\mathcal{R}_{2}^{\prime}}+\mathcal{R}_{1}^{\prime}$
$=$ $~{}B\times(\varepsilon^{4}\partial^{2}\mathcal{T}^{2}D_{t}(\rho D_{t}u))+B\times\varepsilon^{4}\partial^{2}\mathcal{T}^{2}D_{t}(B\times(\nabla\times B))+\mathcal{R}_{1}^{\prime}+\mathcal{R}_{2}^{\prime}$
$=$ $~{}B\times(\varepsilon^{4}\rho\partial^{2}\mathcal{T}^{2}D_{t}^{2}u)+\underbrace{\varepsilon^{4}B\times D_{t}\left(B\times\partial^{2}\mathcal{T}^{2}(\nabla\times B)\right)}_{\mathcal{K}_{1}^{\prime}}+\mathcal{R}_{1}^{\prime}+\mathcal{R}_{2}^{\prime}$
$\underbrace{+B\times\varepsilon^{4}([\partial^{2}\mathcal{T}^{2}D_{t},\rho]D_{t}u)+\varepsilon^{2}B\times\left([\partial^{2}\mathcal{T}^{2},D_{t}](\nabla\times B)+D_{t}\left([\partial^{2}\mathcal{T}^{2},B\times](\nabla\times B)\right)\right)}_{\mathcal{R}_{3}^{\prime}},$ (3.42)
@@
$\int_{\Omega}\left(\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\right)\cdot\left(\varepsilon^{4}B\times D_{t}\left(B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right)\right)\,\mathrm{d}x$
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{3}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x+\frac{1}{2}\int_{\Omega}(\nabla\cdot u)\left|\varepsilon^{3}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x$
$\lesssim$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{3}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x+P(E_{4}(t))E_{5}(t),$ (3.44)
@@
$\mathcal{K}^{\prime}+\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{3}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t))\|\varepsilon^{4}\mathcal{T}^{2}D_{t}^{2}u\|_{2}+P(E_{4}(t),E_{5}(t)).$ (3.45)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}u\right\|_{2}^{2}+\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}B\right\|_{2}^{2}+\left\|\varepsilon^{3}B\times(\nabla\times\mathcal{T}^{2}B)\right\|_{2}^{2}\right)$ (3.46)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}u\right\|_{2}^{2}+\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}B\right\|_{2}^{2}+\left\|\varepsilon^{3}B\times(\nabla\times\mathcal{T}^{2}B)\right\|_{2}^{2}\right)$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}u\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}+1}B\times(\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B)\right\|_{2-k}^{2}\right)$
@@
$\sum_{k=0,1}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{4}\nabla\times\partial_{t}^{k}\mathcal{T}^{4}u\right\|_{1-k}^{2}+\left\|\varepsilon^{4}\nabla\times\partial_{t}^{k}\mathcal{T}^{4}B\right\|_{1-k}^{2}+\left\|\varepsilon^{5}B\times(\nabla\times\partial_{t}^{k}\mathcal{T}^{4}B)\right\|_{1-k}^{2}\right)$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}$ $\left(\left\|\varepsilon^{6}\nabla\times\mathcal{T}^{6}u\right\|_{0}^{2}+\left\|\varepsilon^{6}\nabla\times\mathcal{T}^{6}B\right\|_{0}^{2}+\left\|\varepsilon^{7}B\times(\nabla\times\mathcal{T}^{6}B)\right\|_{0}^{2}\right)\leq P(E_{4}(t)E_{5}(t),E_{6}(t),E_{7}(t))\|\varepsilon^{8}\mathcal{T}^{8}u\|_{0}.$ (3.57)
@@
$\|\varepsilon^{4}\partial_{t}^{k}\nabla\cdot\mathcal{T}^{4}u\|_{1-k}^{2}=$ $\left\|\varepsilon^{6}\partial_{t}^{k}\mathcal{T}^{4}(D_{t}p)\right\|_{1-k}^{2}+\|\varepsilon^{4}[\mathcal{T}^{4},\nabla\cdot]\partial_{t}^{k}u\|_{1-k}^{2},~{}~{}~{}k=0,1,$
$\|\varepsilon^{6}\nabla\cdot\mathcal{T}^{6}u\|_{0}^{2}=$ $\left\|\varepsilon^{8}\mathcal{T}^{6}(D_{t}p)\right\|_{0}^{2}+\|\varepsilon^{6}[\mathcal{T}^{6},\nabla\cdot]u\|_{0}^{2},$ (3.58)
@@
$\sum_{k=0,1}\|\varepsilon^{4}[\mathcal{T}^{4},\nabla\cdot]\partial_{t}^{k}u\|_{1-k}^{2}\leq$ $~{}\varepsilon^{4}E_{5}(t)\lesssim\varepsilon^{4}E_{5}(0)+\int_{0}^{t}E_{6}(\tau)\mathrm{d}\tau,$ (3.59)
$\|\varepsilon^{6}[\mathcal{T}^{6},\nabla\cdot]u\|_{0}^{2}\leq$ $~{}\varepsilon^{4}E_{6}(t)\lesssim\varepsilon^{4}E_{6}(0)+\int_{0}^{t}E_{7}(\tau)\mathrm{d}\tau.$ (3.60)

### v6

@@ 3.2.2 Div-Curl analysis
The term $\|\varepsilon^{4+(k-1)_{+}}\partial_{t}^{k}\mathcal{T}^{2}(D_{t}p)\|_{2-k}^{2}$ will be further reduced to the tangenital estimates of $(u,B)$ by using the argument in Section 3.2.1. Finally, we need to control the tangential derivatives of $u$ and $B$ (including time derivatives) in $E_{5}(t)$ and the $L^{2}$ norm of $\partial_{t}^{5}p$ with suitable Mach number weights.
Next, we analyze the vorticity term. The proof is parallel to Lemma 3.2.
@@
It holds that
$\sum_{k=0}^{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}u\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times\left(\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B\right)\right\|_{2-k}^{2}\right)$ (3.30)
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho\left|\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}u\right|^{2}+\left|\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\right|^{2}+\left|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times(\partial^{2}\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t),E_{5}(t),E_{6}(t)).$
@@
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}|\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B|^{2}\,\mathrm{d}x\underbrace{+\int_{\Omega}\left(\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\right)\cdot\left(\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\nabla\times\mathcal{T}^{2}(BD_{t}p)\right)\,\mathrm{d}x}_{:=\mathcal{K}^{\prime}}$
$\underbrace{-\int_{\Omega}\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\cdot\varepsilon^{2}\left([\partial^{2}\nabla\times\mathcal{T}^{2},D_{t}]B+(u\cdot\nabla)(\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B)+[B\cdot\nabla,\partial^{2}\nabla\times\mathcal{T}^{2}]u\right)\,\mathrm{d}x}_{:=\mathcal{I}_{3}^{\prime}}.$ (3.33)
@@
$\varepsilon^{2}\|\mathcal{T}(\rho u_{j})\partial^{3}\mathcal{T}\partial_{j}u_{i}\|_{0}\leq$ $~{}\|\mathcal{T}(\rho u)\|_{\infty}\|\varepsilon^{2}\mathcal{T}\partial_{j}u\|_{3}\leq P(E_{4}(t),E_{5}(t)),~{}j=1,2.$ (3.37)
@@
For $\mathcal{K}^{\prime}$, we consider the estimate of $\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\nabla\times\mathcal{T}^{2}(BD_{t}p)$. Since,
$\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\nabla\times\mathcal{T}^{2}(BD_{t}p)$
$=$ $~{}\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\mathcal{T}^{2}(D_{t}\nabla p\times B)\underbrace{+\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\mathcal{T}^{2}(\nabla u\cdot\nabla p\times B)+\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\mathcal{T}^{2}(D_{t}p\nabla\times B)}_{:=\mathcal{R}_{1}^{\prime}}$
$=$ $-B\times\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\mathcal{T}^{2}D_{t}\nabla p\underbrace{-\varepsilon^{2}{\mathcal{F}_{p}}[\partial^{2}\mathcal{T}^{2},B\times](D_{t}\nabla p)}_{\mathcal{R}_{2}^{\prime}}+\mathcal{R}_{1}^{\prime}$
$=$ $~{}B\times(\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\mathcal{T}^{2}D_{t}(\rho D_{t}u))+B\times\varepsilon^{2}{\mathcal{F}_{p}}\partial^{2}\mathcal{T}^{2}D_{t}(B\times(\nabla\times B))+\mathcal{R}_{1}^{\prime}+\mathcal{R}_{2}^{\prime}$
$=$ $~{}B\times(\varepsilon^{2}{\mathcal{F}_{p}}\rho\partial^{2}\mathcal{T}^{2}D_{t}^{2}u)+\underbrace{\varepsilon^{2}{\mathcal{F}_{p}}B\times D_{t}\left(B\times\partial^{2}\mathcal{T}^{2}(\nabla\times B)\right)}_{\mathcal{K}_{1}^{\prime}}+\mathcal{R}_{1}^{\prime}+\mathcal{R}_{2}^{\prime}$
$\underbrace{+B\times\varepsilon^{2}{\mathcal{F}_{p}}([\partial^{2}\mathcal{T}^{2}D_{t},\rho]D_{t}u)+\varepsilon^{2}B\times\left([\partial^{2}\mathcal{T}^{2},D_{t}](\nabla\times B)+D_{t}\left([\partial^{2}\mathcal{T}^{2},B\times](\nabla\times B)\right)\right)}_{\mathcal{R}_{3}^{\prime}},$ (3.43)
@@
$\int_{\Omega}\left(\varepsilon^{2}\partial^{2}\nabla\times\mathcal{T}^{2}B\right)\cdot\left(\varepsilon^{2}{\mathcal{F}_{p}}B\times D_{t}\left(B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right)\right)\,\mathrm{d}x$
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x+\frac{1}{2}\int_{\Omega}({D_{t}\mathcal{F}_{p}}+\nabla\cdot u)\left|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x$
$\lesssim$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x+P(E_{4}(t))E_{5}(t),$ (3.45)
@@
$\mathcal{K}^{\prime}+\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times\partial^{2}(\nabla\times\mathcal{T}^{2}B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t))\|\varepsilon^{4}\mathcal{T}^{2}D_{t}^{2}u\|_{2}+P(E_{4}(t),E_{5}(t)).$ (3.46)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}u\right\|_{2}^{2}+\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}B\right\|_{2}^{2}+\left\|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times(\nabla\times\mathcal{T}^{2}B)\right\|_{2}^{2}\right)$ (3.47)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}u\right\|_{2}^{2}+\left\|\varepsilon^{2}\nabla\times\mathcal{T}^{2}B\right\|_{2}^{2}+\left\|\varepsilon^{2}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times(\nabla\times\mathcal{T}^{2}B)\right\|_{2}^{2}\right)$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}u\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times(\nabla\times\mathcal{T}^{2}\partial_{t}^{k}B)\right\|_{2-k}^{2}\right)$
@@
$\sum_{k=0,1}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{4}\nabla\times\partial_{t}^{k}\mathcal{T}^{4}u\right\|_{1-k}^{2}+\left\|\varepsilon^{4}\nabla\times\partial_{t}^{k}\mathcal{T}^{4}B\right\|_{1-k}^{2}+\left\|\varepsilon^{4}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times(\nabla\times\partial_{t}^{k}\mathcal{T}^{4}B)\right\|_{1-k}^{2}\right)$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{6}\nabla\times\mathcal{T}^{6}u\right\|_{0}^{2}+\left\|\varepsilon^{6}\nabla\times\mathcal{T}^{6}B\right\|_{0}^{2}+\left\|\varepsilon^{6}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times(\nabla\times\mathcal{T}^{6}B)\right\|_{0}^{2}\right)$
$\leq$ $~{}P(E_{4}(t),E_{5}(t),E_{6}(t),E_{7}(t))\|\varepsilon^{8}\mathcal{T}^{8}u\|_{0}.$ (3.58)
@@
$\|\varepsilon^{4}\partial_{t}^{k}\nabla\cdot\mathcal{T}^{4}u\|_{1-k}^{2}{\lesssim}$ $\left\|\varepsilon^{6}\partial_{t}^{k}\mathcal{T}^{4}(D_{t}p)\right\|_{1-k}^{2}+\|\varepsilon^{4}[\mathcal{T}^{4},\nabla\cdot]\partial_{t}^{k}u\|_{1-k}^{2},~{}~{}~{}k=0,1,$
$\|\varepsilon^{6}\nabla\cdot\mathcal{T}^{6}u\|_{0}^{2}{\lesssim}$ $\left\|\varepsilon^{8}\mathcal{T}^{6}(D_{t}p)\right\|_{0}^{2}+\|\varepsilon^{6}[\mathcal{T}^{6},\nabla\cdot]u\|_{0}^{2},$ (3.59)
@@
$\sum_{k=0,1}\|\varepsilon^{4}[\mathcal{T}^{4},\nabla\cdot]\partial_{t}^{k}u\|_{1-k}^{2}{\lesssim}$ $~{}\varepsilon^{4}E_{5}(t)\lesssim\varepsilon^{4}E_{5}(0)+\int_{0}^{t}E_{6}(\tau)\mathrm{d}\tau,$ (3.60)
$\|\varepsilon^{6}[\mathcal{T}^{6},\nabla\cdot]u\|_{0}^{2}{\lesssim}$ $~{}\varepsilon^{4}E_{6}(t)\lesssim\varepsilon^{4}E_{6}(0)+\int_{0}^{t}E_{7}(\tau)\mathrm{d}\tau.$ (3.61)